Experiment: Add the 1M NaOH solution, distilled water to a 250mL volumetric flask, what volume of NaOH is needed to prepare .300M of the solution
Answer #1

use dilution formula
M1*V1 = M2*V2
1---> is for stock solution
2---> is for diluted solution

Given:
M1 = 1 M
M2 = 0.3 M
V2 = 250 mL

use:
M1*V1 = M2*V2
V1 = (M2 * V2) / M1
V1 = (0.3*250)/1
V1 = 75 mL
Answer: 75 mL